# Towards Resilient Safety-driven Unlearning for Diffusion Models against Downstream Fine-tuning

### v1

@@ Abstract
Text-to-image (T2I) diffusion models have achieved impressive image generation quality and are increasingly fine-tuned for personalized applications. However, these models often inherit unsafe behaviors from toxic pretraining data, raising growing safety concerns. While recent safety-driven unlearning methods have made promising progress in suppressing model toxicity, they are identified to be fragile to downstream fine-tuning, where we reveal that state-of-the-art methods largely fail to retain their effectiveness even when fine-tuned on entirely benign datasets. To mitigate this problem, in this paper, we propose ResAlign, a safety-driven unlearning framework with enhanced resilience against downstream fine-tuning. By modeling downstream fine-tuning as an implicit optimization problem with a Moreau Envelope-based reformulation, ResAlign enables efficient gradient estimation to minimize the recovery of harmful behaviors. Additionally, a meta-learning strategy is proposed to simulate a diverse distribution of fine-tuning scenarios to improve generalization. Extensive experiments across a wide range of datasets, fine-tuning methods, and configurations demonstrate that ResAlign consistently outperforms prior unlearning approaches in retaining safety after downstream fine-tuning while preserving benign generation capability well.
[…]
@@ 1 Introduction
Text-to-image (T2I) diffusion models have emerged as a dominant class of generative AI due to their unprecedented ability to synthesize high-quality, diverse, and aesthetically compelling general images from natural language descriptions rombach2022ldm ; podell2024sdxl . Beyond synthesizing general images, there is also a growing interest in customizing pretrained models for personalized generation, e.g., generating images of specific facial identities or artistic styles that are underrepresented in the original training data. This is typically achieved by fine-tuning the pretrained base model on a small reference dataset for a few steps ruiz2023dreambooth ; li2024towards ; wang2025sleepermark . The development of several advanced fine-tuning methods hu2022lora ; ruiz2023dreambooth ; han2023svdiff as well as the rapid proliferation of “fine-tuning-as-a-service” platforms civitai2025 has made personalization widely accessible, fueling a surge in applications such as stylized avatars, fan art, and thematic illustrations, which are becoming increasingly popular especially among younger users wang2023factors ; commonsense2024teens ; ali2023constructing .
Yet alongside the rapid advancements of diffusion models, growing concerns have emerged regarding their potential to generate inappropriate or harmful content (e.g., sexually explicit imagery) gandikota2023erasing ; schramowski2023sld ; li2025rethinking . Due to the large-scale and web-crawled nature of their training datasets, modern T2I models inevitably ingest amounts of harmful materials during pretraining schuhmann2022laion . As a result, these models can reproduce such content either when explicitly prompted or inadvertently triggered. For example, recent studies schramowski2023sld ; qu2023unsafe ; li2024art based on real-world user generations demonstrate that widely-deployed models like Stable Diffusion rombach2022ldm are particularly prone to producing unsafe content, even though many of the prompts that lead to unsafe outputs appear benign and may not be intended to generate harmful results. These vulnerabilities not only allow malicious exploitation through direct or adversarial prompting yang2024sneakyprompt ; zhang2024unlearndiff , but also increase the risk of harmful unintended exposure for ordinary, benign users li2024art , raising serious ethical concerns for real-world deployment. In response to these concerns, a variety of safety-driven unlearning methods gandikota2023erasing ; li2024safegen ; zhang2024defensive have recently been proposed to modify the pretrained models’ parameters in order to suppress their capacity for unsafe generation.
@@
In conclusion, our main contributions are threefold. (1) We empirically reveal that existing safety-driven unlearning methods largely fail to retain their effectiveness after downstream fine-tuning, even when the data does not contain unsafe content. (2) We propose ResAlign, a resilient safety-driven unlearning framework for T2I diffusion models. By leveraging a Moreau Envelope-based approximation and a meta-learning strategy over diverse adaptation scenarios, ResAlign explicitly accounts for and minimizes post-unlearning degradation due to downstream fine-tuning efficiently with high generalizability. We further provide theoretical insights to help understand the empirical effectiveness of our method. (3) Through extensive experiments, we show that ResAlign consistently outperforms baselines in maintaining safety after fine-tuning, and generalizes well to a wide range of advanced fine-tuning methods, datasets, and hyperparameters. It also preserves both general and personalized generation quality well, and remains effective to multiple diffusion architectures, unsafe concepts, and even when the data is harmfully contaminated or adaptively attacked.
@@ 2 Background & Related Work
Mitigating Unsafe Generation in T2I Models. To mitigate unsafe generation in large models, several strategies have been developed recently qu2024provably ; yang2024guardti ; liu2024latent ; gandikota2023erasing , which can be broadly categorized into detection-based and unlearning-based. Detection-based strategies deploy external safety filters (e.g., DNN-based harmful prompt/image detectors) to inspect and block unsafe requests compvis2022sc ; liu2024latent . While effective, it does not inherently “detoxify” the diffusion model, often introducing non-negligible additional computational and memory overhead during inference li2024safegen , and can be easily removed when the model is opensourced fernandez2023stable ; rando2022red . On the other hand, unlearning-based approaches internalize safety constraints by directly modifying model parameters to reduce unsafe generation at its source gandikota2023erasing ; kumari2023ablating ; li2024safegen . From a high level, given a pretrained T2I diffusion model $\theta\in\mathbb{R}^{d}$, existing safety-driven unlearning approaches typically aim to optimize the model with the following objective:
@@
where $\mathcal{L}_{\text{harmful}}(\theta)$ is a harmful loss such that minimizing it encourages the model to unlearn harmful behaviors, and $\mathcal{R}(\theta)$ is an utility-preserving regularization term that maintains the model’s benign generation capabilities. The hyperparameter $\alpha>0$ is a weighting parameter. Since the seminal work of CA kumari2023ablating and ESD gandikota2023erasing , safety-driven unlearning has rapidly gained traction, with a growing line of subsequent works expanding the design space across multiple dimensions including loss function design kumari2023ablating ; gandikota2023erasing , set of updated parameters li2024safegen ; gandikota2023erasing , optimization strategies gandikota2024unified ; zhang2024defensive ; gong2024rece , etc.
Despite these progresses, a recent pioneering work pan2024leveraging has identified a malicious fine-tuning issue, where unlearned models can quickly regain harmful generative abilities after fine-tuning on a small set of unsafe images, producing diverse outputs that resemble the quality and variety of the original pre-unlearned model. To the best of our knowledge, the most advanced attempt to address this issue is LCFDSD pan2024leveraging , which encourages separation in the latent space between clean and harmful data distributions to increase the difficulty of recovering unsafe behaviors. However, this comes with a notable degradation on the benign utility of the model and only brings slight improvements over existing unlearning methods, as validated in our experiments (Sec. 4.2). To summarize, how to effectively mitigate the post-finetuning resurgence of harmful capabilities in unlearned models remains an important yet largely unaddressed problem and is worth further exploration.
@@ 3 Methodology
Motivation & Problem Formulation. Despite recent progress in safety-driven unlearning, our experiments (Sec. 4.2) reveal that many state-of-the-art methods fail to retain their effectiveness after downstream fine-tuning, even when the adaptation data is entirely benign (instead of solely with harmful data). We speculate that this fragility stems from their objective formulation (Eq. (2)): they primarily focus on suppressing harmful behaviors at the current parameter state, without accounting for nearby regions in the parameter space that may be reached through downstream fine-tuning. As a result, the local neighborhood of the unlearned model may still be vulnerable, such that even benign gradient signals during fine-tuning can inadvertently push the model into toxic regions and erode its safety. Motivated by this observation, we aim to learn model parameters that are not only safe in their current state, but also lie in regions of the parameter space where safety is preserved under downstream fine-tuning. In other words, we seek to explicitly penalize the increase in harmfulness loss caused by (simulated) downstream fine-tuning during unlearning. Formally, given a pre-trained T2I diffusion model parameterized by $\theta\in\mathbb{R}^{d}$, we propose the following resilient unlearning objective:
@@
where $\theta_{\text{FT}}^{*}=\textsc{Adapt}(\theta,\mathcal{D}_{\text{FT}},\mathcal{C})$ represents the parameters obtained by adapting $\theta$ on dataset $\mathcal{D}_{\text{FT}}$ with configuration $\mathcal{C}$. For example, if $\mathcal{C}$ is the standard gradient descent for $T$ steps with a learning rate of $\eta$, then we have $\theta_{\text{FT}}^{*}=\theta-\eta\sum_{t=0}^{T-1}\nabla_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$, where $\mathcal{L}_{\text{FT}}$ represents the standard diffusion denoising loss and $\theta^{(i)}$ indicates the model parameters at the $i$-th step with $\theta^{(0)}=\theta$. $\alpha$ and $\beta$ are hyperparameters that balance the three terms. In this paper, we follow previous works zheng2024imma ; zhang2024defensive and instantiate $\mathcal{L}_{\text{harmful}}(\theta)$ as the negative denoising score matching loss on a set of inappropriate prompt-image pairs and $\mathcal{R}(\theta)$ as the distillation loss that minimizes the difference of noise prediction between the current model and the original model $\hat{\epsilon}_{\theta_{0}}$ on a set of preservation prompts, i.e., $\mathcal{L}_{\text{harmful}}(\theta)=-\mathbb{E}_{\epsilon,t,(x,p)\sim\mathcal{D}_{\text{harmful}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(\alpha_{t}x+\sigma_{t}\epsilon,p,t)-\epsilon\right\|_{2}^{2}]$ and $\mathcal{R}(\theta)=\mathbb{E}_{x_{T}\sim\mathcal{N}(0,1),t,p\sim\mathcal{P}_{\text{preserve}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(x_{t},p,t)-\hat{\epsilon}_{\theta_{0}}(x_{t},p,t)\right\|_{2}^{2}]$. Note that our method is also compatible with other designs of these losses, as long as they offer similar utility.
@@
The first two terms in Eq. (4) are standard and can be easily computed via backpropagation. The difficulty lies in the last term, i.e., the hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$. Because $\theta^{*}_{\text{FT}}$ is an implicit function of $\theta$, according to the chain rule, we have $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})=\left(\partial\theta^{*}_{\text{FT}}/\partial\theta\right)^{\top}\cdot\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$, which involves the Jacobian ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}$. Since $\theta^{*}_{\text{FT}}$ is obtained through an optimization procedure (e.g., $T$-step gradient descent), this derivative expands into a product of Jacobians ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}=\prod_{t=0}^{T-1}\left(I-\eta\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})\right)$, where $\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$ is the Hessian matrix of the fine-tuning loss w.r.t. $\theta^{(t)}$. As such, direct computation requires storing and backpropagating through the entire trajectory and their Hessian matrices, which is computationally and memory-prohibitive.
To address this issue, inspired by recent work in bilevel optimization and adversarial training liu2024moreau ; gao2023moreau ; zhang2024stability ; rajeswaran2019meta ; xiao2024uniformly , we approximate fine-tuning as an implicit optimization over the Moreau Envelope moreau1965proximite of the loss. Specifically, we approximately regard the solution of downstream fine-tuning as the minimizer of the following Moreau Envelope (ME) rajeswaran2019meta , that is,
@@
where $\gamma>0$ is a proximity coefficient rajeswaran2019meta . Intuitively, in the context of diffusion model fine-tuning, we typically start from a strong pretrained base model and only take a few gradient steps. This implicitly restricts the solution $\theta_{\text{FT}}^{*}$ to remain near $\theta$, making it well-approximated by the minimizer of the proximal objective in Eq. (5). As a result, $\theta_{\text{FT}}^{*}$ naturally satisfies the first-order optimality condition, i.e., $\nabla_{\theta^{*}_{\text{FT}}(\theta)}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}(\theta_{\text{FT}}^{*}-\theta)=0$ according to the Karush-Kuhn-Tucker theorem kuhn2013nonlinear ; rajeswaran2019meta . Since $\theta_{\text{FT}}^{*}$ is implicitly defined as a function of $\theta$ through this optimality condition, we apply the implicit function theorem to differentiate both sides with respect to $\theta$. Then, by right-multiplying both sides by $\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, we can derive the following equation (details in Appendix A):
@@
Let $A:=\nabla^{2}_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{FT}}+\frac{1}{\gamma}I$, $x:=\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, and $b:=\frac{1}{\gamma}\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}$, we observe from Eq. (6) that the hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$ is essentially the solution to the linear system $Ax=b$, which can be efficiently solved using the Richardson iteration method richardson1911ix :
$x^{(k+1)}=b-\frac{1}{\gamma}\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta^{*}_{\text{FT}},\mathcal{D}_{\text{FT}})\cdot x^{(k)},$ (7)
@@
Our approach has the following advantages. First, computing the implicit gradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ only depends on the final fine-tuned parameters $\theta_{\text{FT}}^{*}$ and the local Hessian-vector product $\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}\cdot x$. This product can be efficiently computed as a Hessian-vector product (HVP), which is readily supported in modern autodiff frameworks such as PyTorch (via backward-mode Auto Differentiation), without explicitly forming or inverting any second-order Hessian matrices.
@@
3:$x^{(0)}\leftarrow 0$, $b\leftarrow\frac{1}{\gamma}\nabla_{\theta_{\text{FT}}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}})$
4:for $k=0$ to $K-1$ do
@@
7:    $x^{(k+1)}\leftarrow b-\frac{1}{\gamma}\cdot\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}})\cdot x^{(k)}$
@@
Besides, as the gradient is entirely determined by $\theta_{\text{FT}}^{*}$, there is no need to store the intermediate models during fine-tuning, which significantly reduces memory and computational overhead. Finally, we observe that the Richardson iteration used to approximate the implicit gradient (i.e., Eq. (7)) converges quickly within a few steps (5 in this paper), enabling efficient training for large-scale diffusion models. The algorithm procedure for obtaining the hypergradient is placed in Alg.1.
@@
1:Input: Initial model parameters $\theta_{0}$, number of outer loop iterations $I$, number of inner loop iterations $J$, dataset $\mathcal{D}$, distribution over configurations $\pi(\mathcal{C})$, loss functions $\mathcal{L}_{\text{FT}},\mathcal{L}_{\text{harmful}},\mathcal{R}$, learning rate $\eta$, hyperparameters $\alpha,\beta,\gamma,K$
@@
3:for $i=0$ to $I-1$ do $\triangleright$ Outer Loop
@@
10:    $g_{i}\leftarrow g_{i}+\frac{\beta}{J}\sum^{J}_{j}g_{i,j}$triangleright Aggregating hypergradients
@@
In general, Proposition 1 shows that our additional term can be regarded as an implicit penalty on the trace of the Hessian of the harmfulness loss with respect to $\theta$. The trace of the Hessian is a well-known indicator of the overall curvature of the loss landscape hutchinson1989stochastic ; foret2021sharpnessaware , where larger values reflect sharper minima that are highly sensitive to parameter perturbations, while smaller values correspond to flatter and more stable regions. Interestingly, prior works in generalization and robustness keskar2016large ; foret2021sharpnessaware have shown that SGD-style optimizers tend to lead models to sharp minima, i.e., regions in the loss landscape with high curvature that are sensitive to parameter perturbations. This connection offers a potential explanation for the fragility of existing unlearning methods: since they do not explicitly regularize the curvature of the harmfulness loss, the resulting models, while appearing safe at their current parameter state, may lie in locally sharp and unstable regions of the loss landscape. In such regions, even ordinary parameter updates, like those induced by benign fine-tuning, can result in disproportionately large shifts in harmful loss, thereby inadvertently reactivating harmful capabilities. In contrast, our ResAlign introduces an implicit penalty on the trace of the Hessian of the harmfulness loss, encouraging convergence to flatter regions of the loss surface. This may help reduce the model’s sensitivity to downstream updates and thus improves its resilience to post-unlearning fine-tuning. We provide a proof of Proposition 1 and further discussion in Appendix A.

### v2

@@ Abstract
Text-to-image (T2I) diffusion models have achieved impressive image generation quality and are increasingly fine-tuned for personalized applications. However, these models often inherit unsafe behaviors from toxic pretraining data, raising growing safety concerns. While recent safety-driven unlearning methods have made promising progress in suppressing model toxicity, they are found to be fragile to downstream fine-tuning, as we reveal that state-of-the-art methods largely fail to retain their effectiveness even when fine-tuned on entirely benign datasets. To mitigate this problem, in this paper, we propose ResAlign, a safety-driven unlearning framework with enhanced resilience against downstream fine-tuning. By modeling downstream fine-tuning as an implicit optimization problem with a Moreau envelope-based reformulation, ResAlign enables efficient gradient estimation to minimize the recovery of harmful behaviors. Additionally, a meta-learning strategy is proposed to simulate a diverse distribution of fine-tuning scenarios to improve generalization. Extensive experiments across a wide range of datasets, fine-tuning methods, and configurations demonstrate that ResAlign consistently outperforms prior unlearning approaches in retaining safety, while effectively preserving benign generation capability. Our code and pretrained models are publicly available here.
. Disclaimer: This paper includes AI-generated images containing partially nude human figures and other sensitive content, shown only for research purposes.
@@ 1 Introduction
Text-to-image (T2I) diffusion models have emerged as a dominant class of generative AI due to their unprecedented ability to synthesize high-quality, diverse, and aesthetically compelling general images from natural language descriptions rombach2022ldm; podell2024sdxl. Beyond synthesizing general images, there is also a growing interest in customizing pretrained models for personalized generation, e.g., generating images of specific facial identities or artistic styles that are underrepresented in the original training data li2025rethinking. This is typically achieved by fine-tuning the pretrained base model on a small reference dataset for a few steps ruiz2023dreambooth; li2024towards; wang2025sleepermark. The development of several advanced fine-tuning methods hu2022lora; ruiz2023dreambooth; han2023svdiff as well as the rapid proliferation of “fine-tuning-as-a-service” platforms civitai2025 has made personalization widely accessible, fueling a surge in applications such as stylized avatars, fan art, and thematic illustrations, which are becoming increasingly popular especially among younger users wang2023factors.
Yet alongside the rapid advancements of diffusion models, growing concerns have emerged regarding their potential to generate inappropriate or harmful content (e.g., sexually explicit imagery) gandikota2023erasing; schramowski2023sld. Due to the large-scale and web-crawled nature of their training datasets, modern T2I models inevitably ingest amounts of harmful material during pretraining schuhmann2022laion. As a result, these models can reproduce such content either when explicitly prompted or inadvertently triggered. For example, recent studies schramowski2023sld; qu2023unsafe; li2024art based on real-world user generations demonstrate that widely-deployed models like Stable Diffusion rombach2022ldm are particularly prone to producing unsafe content, even though many of the prompts that lead to unsafe outputs appear benign and may not be intended to generate harmful results. These vulnerabilities not only allow malicious exploitation through direct or adversarial prompting yang2024sneakyprompt; zhang2024unlearndiff, but also increase the risk of harmful unintended exposure for ordinary, benign users li2024art, raising serious ethical concerns for real-world deployment. In response to these concerns, a variety of safety-driven unlearning methods gandikota2023erasing; li2024safegen; zhang2024defensive have recently been proposed to modify the pretrained models’ parameters in order to suppress their capacity for unsafe generation.
@@
In conclusion, our main contributions are threefold. (1) We empirically reveal that existing safety-driven unlearning methods largely fail to retain their effectiveness after downstream fine-tuning, even when the data does not contain unsafe content. (2) We propose ResAlign, a resilient safety-driven unlearning framework for T2I diffusion models. By leveraging a Moreau envelope-based approximation and a meta-learning strategy over diverse adaptation scenarios, ResAlign explicitly accounts for and minimizes post-unlearning degradation due to downstream fine-tuning efficiently with high generalizability. We further provide theoretical insights to help understand the empirical effectiveness of our method. (3) Through extensive experiments, we show that ResAlign consistently outperforms baselines in maintaining safety after fine-tuning, and generalizes well to a wide range of advanced fine-tuning methods, datasets, and hyperparameters. It also preserves both general and personalized generation quality well, generalizes across various diffusion models and loss functions, and remains effective even under harmful data contamination or adaptive attacks.
@@ 2 Background & Related Work
Mitigating Unsafe Generation in T2I Models. To mitigate unsafe generation in T2I models, several strategies have been developed recently qu2024provably; yang2024guardti; liu2024latent; gandikota2023erasing, which can be broadly categorized into detection-based and unlearning-based. Detection-based strategies deploy external safety filters (e.g., DNN-based harmful prompt/image detectors) to inspect and block unsafe requests compvis2022sc; liu2024latent. While effective, they do not inherently “detoxify” the diffusion model, often introducing non-negligible additional computational and memory overhead during inference li2024safegen, and can be easily removed once the model is open-sourced rando2022red. On the other hand, unlearning-based approaches internalize safety constraints by directly modifying model parameters to reduce unsafe generation at its source gandikota2023erasing; kumari2023ablating; li2024safegen. At a high level, given a pretrained T2I diffusion model $\theta\in\mathbb{R}^{d}$, existing safety-driven unlearning approaches typically aim to optimize the model with the following objective:
@@
where $\mathcal{L}_{\text{harmful}}(\theta)$ is a harmful loss such that minimizing it encourages the model to unlearn harmful behaviors, and $\mathcal{R}(\theta)$ is a utility-preserving regularization term that maintains the model’s benign generation capabilities. The hyperparameter $\alpha>0$ is a weighting coefficient. Since the seminal work of CA kumari2023ablating and ESD gandikota2023erasing, safety-driven unlearning has rapidly gained traction, with a growing line of subsequent works expanding the design space across multiple dimensions including loss function design kumari2023ablating; gandikota2023erasing, sets of updated parameters li2024safegen; gandikota2023erasing, optimization strategies gandikota2024unified; zhang2024defensive; gong2024rece, and others.
Despite this progress, a recent pioneering work pan2024leveraging has identified a malicious fine-tuning issue, where unlearned models can quickly regain harmful generative abilities after fine-tuning on a small set of unsafe images, producing diverse outputs that resemble the quality and variety of the original pre-unlearned model. To the best of our knowledge, the most advanced attempt to address this issue is LCFDSD pan2024leveraging, which encourages separation in the latent space between clean and harmful data distributions to increase the difficulty of recovering unsafe behaviors. However, this comes with a notable degradation in the benign utility of the model and only brings slight improvements over existing unlearning methods, as validated in our experiments (Sec. 4.2). To summarize, how to effectively mitigate the post-fine-tuning resurgence of harmful capabilities in unlearned models remains an important yet largely unaddressed problem and is worth further exploration.
@@ 3 Methodology
Motivation & Problem Formulation. Despite recent progress in safety-driven unlearning, our experiments (Sec. 4.2) reveal that many state-of-the-art methods fail to retain their effectiveness after downstream fine-tuning, even when the adaptation data is entirely benign (instead of solely with harmful data). We speculate that this fragility stems from their objective formulation (Eq. (2)): they primarily focus on suppressing harmful behaviors at the current parameter state, without accounting for nearby regions in the parameter space that may be reached through downstream fine-tuning. As a result, the local neighborhood of the unlearned model may still be vulnerable, such that even benign gradient signals during fine-tuning can inadvertently push the model into toxic regions and erode its safety. Motivated by this observation, we aim to learn model parameters that are not only safe in their current state, but also lie in regions of the parameter space where safety is preserved under downstream fine-tuning. In other words, we seek to explicitly penalize the increase in harmfulness loss caused by (simulated) downstream fine-tuning during unlearning. Formally, given a pretrained T2I diffusion model parameterized by $\theta\in\mathbb{R}^{d}$, we propose the following resilient unlearning objective:
@@
where $\theta_{\text{FT}}^{*}=\textsc{Adapt}(\theta,\mathcal{D}_{\text{FT}},\mathcal{C})$ represents the parameters obtained by adapting $\theta$ on dataset $\mathcal{D}_{\text{FT}}$ with configuration $\mathcal{C}$. For example, if $\mathcal{C}$ is the standard gradient descent for $T$ steps with a learning rate of $\eta$, then we have $\theta_{\text{FT}}^{*}=\theta-\eta\sum_{t=0}^{T-1}\nabla_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$, where $\mathcal{L}_{\text{FT}}$ represents the standard diffusion denoising loss and $\theta^{(i)}$ indicates the model parameters at the $i$-th step with $\theta^{(0)}=\theta$. $\alpha$ and $\beta$ are hyperparameters that balance the three terms. In this paper, we follow previous works zheng2024imma; zhang2024defensive and instantiate $\mathcal{L}_{\text{harmful}}(\theta)$ as the negative denoising score matching loss on a set of inappropriate prompt-image pairs and $\mathcal{R}(\theta)$ as the distillation loss that minimizes the difference of noise prediction between the current model and the original model $\hat{\epsilon}_{\theta_{0}}$ on a set of preservation prompts, i.e., $\mathcal{L}_{\text{harmful}}(\theta)=-\mathbb{E}_{\epsilon,t,(x,p)\sim\mathcal{D}_{\text{harmful}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(\alpha_{t}x+\sigma_{t}\epsilon,p,t)-\epsilon\right\|_{2}^{2}]$ and $\mathcal{R}(\theta)=\mathbb{E}_{\epsilon,t,(x_{0},p)\sim\mathcal{D}_{\text{preserve}}}[w_{t}\cdot\left\|\hat{\epsilon}_{\theta}(x_{t},p,t)-\hat{\epsilon}_{\theta_{0}}(x_{t},p,t)\right\|_{2}^{2}]$. Note that our method is also compatible with other designs of these losses, as long as they offer similar utility (see results in Sec. 5).
@@
The first two terms in Eq. (4) are standard and can be easily computed via backpropagation. The difficulty lies in the last term, i.e., the hypergradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$. Because $\theta^{*}_{\text{FT}}$ is an implicit function of $\theta$, according to the chain rule, we have $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})=\left(\partial\theta^{*}_{\text{FT}}/\partial\theta\right)^{\top}\cdot\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$, which involves the Jacobian ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}$. Since $\theta^{*}_{\text{FT}}$ is typically obtained through $T$-step iterative gradient descent, this derivative expands into a product of Jacobians ${\partial\theta_{\text{FT}}^{*}}/{\partial\theta}=\prod_{t=0}^{T-1}\left(I-\eta\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})\right)$, where $\nabla^{2}_{\theta^{(t)}}\mathcal{L}_{\text{FT}}(\theta^{(t)},\mathcal{D}_{\text{FT}})$ is the Hessian matrix of the fine-tuning loss w.r.t. $\theta^{(t)}$. As such, direct computation requires storing and backpropagating through the entire trajectory and their Hessian matrices, which is computationally and memory-prohibitive.
To address this issue, inspired by recent work in bilevel optimization liu2024moreau; gao2023moreau; zhang2024stability; rajeswaran2019meta, we approximate fine-tuning as an implicit optimization over the Moreau envelope moreau1965proximite of the loss. Specifically, we approximately regard the solution of downstream fine-tuning (i.e., the resulting parameters $\theta^{*}_{\text{FT}}$) as the minimizer of the following Moreau envelope (ME) moreau1965proximite; gao2023moreau; rajeswaran2019meta, i.e.,
@@
where $\gamma$ (set as $1$ in this paper) is a proximity coefficient rajeswaran2019meta. Intuitively, when fine-tuning diffusion models, we typically start from a strong pretrained base model and only take a few gradient steps. This implicitly restricts the solution $\theta_{\text{FT}}^{*}$ to remain near $\theta$, making it well-approximated by the minimizer of the proximal objective in Eq. (5). As a result, $\theta_{\text{FT}}^{*}$ naturally satisfies the first-order optimality condition, i.e., $\nabla_{\theta^{*}_{\text{FT}}(\theta)}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}}^{*},\mathcal{D}_{\text{FT}})+\frac{1}{\gamma}(\theta_{\text{FT}}^{*}-\theta)=0$ according to the Karush-Kuhn-Tucker theorem kuhn2013nonlinear; rajeswaran2019meta. Since $\theta_{\text{FT}}^{*}$ is implicitly defined as a function of $\theta$ through this optimality condition, we apply the implicit function theorem to differentiate both sides with respect to $\theta$. Then, by right-multiplying both sides by $\nabla_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, we can derive the following equation (details in Appendix A):
@@
Let $A:=\nabla^{2}_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{FT}}+\frac{1}{\gamma}I$, $x:=\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$, and $b:=\frac{1}{\gamma}\nabla_{\theta^{*}_{\text{FT}}}\mathcal{L}_{\text{harmful}}$, we observe from Eq. (6) that $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta^{*}_{\text{FT}})$ is essentially the solution to the linear system $Ax=b$, which can be efficiently solved using the Richardson iteration method richardson1911ix (setting the relaxation parameter as $\omega=\gamma=1$):
$x^{(k+1)}=\gamma b-{\gamma}\nabla^{2}_{\theta_{\text{FT}}^{*}}\mathcal{L}_{\text{FT}}(\theta^{*}_{\text{FT}},\mathcal{D}_{\text{FT}})\cdot x^{(k)},$ (7)
@@
Our approach has the following advantages. First, computing the implicit gradient $\nabla_{\theta}\mathcal{L}_{\text{harmful}}(\theta_{\text{FT}}^{*})$ only depends on the final fine-tuned parameters $\theta_{\text{FT}}^{*}$ and the local Hessian-vector product $\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}\cdot x$. This product can be efficiently computed as a Hessian-vector product (HVP), which is readily supported in modern autodiff frameworks such as PyTorch (via backward-mode automatic differentiation),
@@
[…]
4:for $k=0$ to K1K-1 do
@@
7:  $x^{(k+1)}\leftarrow\gamma b-{\gamma}\nabla^{2}_{\theta_{\text{FT}}}\mathcal{L}_{\text{FT}}(\theta_{\text{FT}})\cdot x^{(k)}$
@@
without explicitly forming or inverting any second-order Hessian matrices. Besides, as the gradient is entirely determined by $\theta_{\text{FT}}^{*}$, there is no need to store the intermediate models during fine-tuning, which significantly reduces memory overhead. Finally, we observe that the Richardson iteration used to approximate the implicit gradient (i.e., Eq. (7)) converges quickly within a few steps (5 in this paper), enabling efficient training for large-scale diffusion models. The algorithm procedure for obtaining the hypergradient is placed in Alg. 1.
@@
1:Input: Initial model parameters $\theta_{0}$, number of outer loop iterations II, number of inner loop iterations $J$, dataset $\mathcal{D}$, distribution over configurations $\pi(\mathcal{C})$, loss functions $\mathcal{L}_{\text{FT}},\mathcal{L}_{\text{harmful}},\mathcal{R}$, learning rate $\eta$, hyperparameters $\alpha,\beta,\gamma,K$
@@
3:for $i=0$ to I1I-1 do $\triangleright$ Outer Loop
@@
10:  $g_{i}\leftarrow g_{i}+\frac{\beta}{J}\sum^{J}_{j}g_{i,j}$ $\triangleright$ Aggregate hypergradients
@@
In general, Proposition 1 shows that our additional term can be regarded as an implicit penalty on the trace of the Hessian of the harmfulness loss with respect to $\theta$. The trace of the Hessian is a well-known indicator of the overall curvature of the loss landscape hutchinson1989stochastic; foret2021sharpnessaware, where larger values reflect sharper minima that are highly sensitive to parameter perturbations, while smaller values correspond to flatter and more stable regions. Interestingly, prior works in generalization and robustness keskar2016large; foret2021sharpnessaware have shown that SGD-style optimizers tend to lead models to sharp minima, i.e., regions in the loss landscape with high curvature that are sensitive to parameter perturbations. This connection offers a potential explanation for the fragility of existing unlearning methods: since they do not explicitly regularize the curvature of the harmfulness loss, the resulting models, while appearing safe at their current parameter state, may lie in locally sharp and unstable regions of the loss landscape. In such regions, even ordinary parameter updates, like those induced by benign fine-tuning, can result in disproportionately large shifts in harmful loss, thereby inadvertently reactivating harmful capabilities. In contrast, our ResAlign introduces an implicit penalty on the trace of the Hessian of the harmfulness loss, encouraging convergence to flatter regions of the loss surface. This may help reduce the model’s sensitivity to downstream updates and thereby improves its resilience to post-unlearning fine-tuning. We provide a proof of Proposition 1 and further discussion in Appendix A.